OpenStudy (anonymous):

The number 20, 21, 29. form a Pythagorean triple. Which other set of numbers is a Pythagorean triple?

OpenStudy (anonymous):

A pythagorean triple is the group of numbers, (a, b, c) where a^2 + b^2 = c^2. So think, what is another triple that can be done. What if, using the equation above, you can your first number to 21 as well? 21^2 + 21^2 = c^2. Solve for c!

OpenStudy (anonymous):

Answer Choices: A. 21, 22, 30 B. 40, 42, 45 C. 60, 66, 87 D. 70, 84, 116